### v1

@@ 3.1. Pencils of quadrics
A vector $\mathrm{v}\in T_{P}\operatorname{Gr}_{d}(\mathcal{Q})$ is fitting if and only if for every $\mathcal{C}^{1}$ curve $\gamma:[0,1]\to\operatorname{Gr}_{d}(\mathcal{Q})$ with $\gamma_{0}=P$ and $\gamma^{\prime}_{0}=\mathrm{v}$, $(\gamma_{t},\gamma_{0})$ is a fitting pair for all $t>0$ small enough.
@@
Let us fix a complement $H$ of $P^{\circ}$ in $S^{2}V$. Let $N=\dim(\mathcal{Q})=\frac{n(n+1)}{2}$. We identify a neighborhood of $P^{\circ}\subset\operatorname{Gr}_{N-d}(\mathcal{Q})$ with $\operatorname{Hom}(P^{\circ},H)$. If there exist a non-zero element $p\in\operatorname{Ker}(\mathrm{v}^{\circ})\cap S^{2}V^{\geq 0}$, one can consider the curve where $\gamma_{t}$ corresponds to
@@
The non-zero element $p\in S^{2}V^{\geq 0}$ belongs to $P^{\circ}=\gamma_{0}^{\circ}$ and $\gamma_{t}^{\circ}$ hence the pair $(\gamma_{0},\gamma_{t})$ is not fitting for any $t>0$.
In general the element corresponding to $\gamma_{t}$ is equal for $t$ close to 00 to the element of $\operatorname{Gr}_{N-d}(\mathcal{Q})$ which is the graph of the map $P^{\circ}\to H$:
@@
In this case the distance between $\mathbb{P}(\gamma_{0}^{\circ}\cap S^{2}V^{\geq 0})$ and $\mathbb{P}(\gamma_{t}^{\circ}\cap S^{2}V^{\geq 0})$ grows at least linearly in $t$ for $t$ close to 00. ∎
@@
As in the third item of Proposition 3.10 a vector is well-fitting if and only if it belongs to $C_{[q]}$ for some $[q]\in\mathbb{S}\mathcal{E}_{P}$. We just check that the sets $C_{[q]}$ are indeed open convex cones.
@@
Consider the tautological rank $d$ vector bundle $p:\mathcal{E}\to\operatorname{Gr}_{d}(\mathcal{Q})$, which admits a tautological projection $\pi:\mathcal{E}\to\mathcal{Q}$.
@@
    for all $v\in T_{x_{0}}M$ there exist $\mathrm{w}\in T\mathcal{E}$ such that $\mathrm{d}p(\mathrm{w})=\mathrm{d}u(v)$ and $\mathrm{d}\pi(\mathrm{w})\in T\mathcal{Q}$ is positive.
Note that that an element in $T\mathcal{Q}$ is a pair $(q,\dot{q})$ and we say that it is positive if the tangent vector $\dot{q}\in\mathcal{Q}$ is positive.
@@
A pair $(P,q_{0})\in\operatorname{Gr}_{d}(\mathcal{Q})\times\mathcal{Q}$ correspond to an element in $\mathcal{E}$ if and only if $p\in P$. A pair $\left(\mathrm{v},\dot{q}\right)\in T_{P}\operatorname{Gr}_{2}(\mathcal{Q})\times T_{q}\mathcal{Q}$ can be written as $\left(\mathrm{d}p(\mathrm{w}),\mathrm{d}\pi(\mathrm{w})\right)$ for some $\mathrm{w}\in T\mathcal{E}$ if and only if $\dot{q}$ belongs to the class defined by $\mathrm{v}(q_{0})$.
@@
The pullback $u^{*}\mathcal{E}$ of the tautological bundle $p:\mathcal{E}\to\operatorname{Gr}_{d}(\mathcal{Q})$ defines a rank $d$ vector bundle over $M$. We define the sphere bundle $\mathbb{S}u^{*}\mathcal{E}$ as the quotient of the vector bundle $u^{*}\mathcal{E}$ by the action of positive scalars.
Recall that $M$ has dimension $d$. We consider flows on $\mathbb{S}u^{*}\mathcal{E}$ so that some form of contraction occurs along the flow lines. We denote also by $p$ the bundle maps $u^{*}\mathcal{E}\to M$, $\mathbb{S}u^{*}\mathcal{E}\to M$, with a slight abuse of notations.
The fiber at $P\in\operatorname{Gr}_{d}(\mathcal{Q})$ of the bundle $\mathcal{E}$ is identified with the vector subspace $P\subset\mathcal{Q}$. Since all the fibers are naturally identified with subsets of $\mathcal{Q}$, there is a natural projection $\pi:\mathcal{E}\to\mathcal{Q}$. We still denote by $\pi:u^{*}\mathcal{E}\to\mathcal{Q}$ the corresponding projection with a slight abuse of notations.
@@
A fitting flow for a continuous map $u:M\to\operatorname{Gr}_{d}(\mathcal{Q})$ is a continuous flow $(\Phi_{t})_{t\in\mathbb{R}}$ on $\mathbb{S}u^{*}\mathcal{E}$ such that one can choose a representative $q^{\prime}$ of $[q^{\prime}]=\Phi_{t}(q)$ such that $\pi(q^{\prime})-\pi(q)\in\mathcal{Q}$ is positive.
@@
Assume by contradiction that for some $t_{0}>0$ and $q\in\mathbb{S}u^{*}\mathcal{E}$ one has $u(x)=p(q)=p(\Phi_{t_{0}}(q))$. The fact that the flow is fitting implies that for some $\lambda>0$, $\lambda\pi(\Phi_{t_{0}}(q))-\pi(q)\in\mathcal{Q}$ is positive. This positive quadric belongs to $u(x)$, contradicting the fact that $u(x)\in\operatorname{Gr}_{d}^{{\operatorname{mix}}}(\mathcal{Q})$.
@@
Let $u:M\to\operatorname{Gr}^{{\operatorname{mix}}}_{d}(\mathcal{Q})$ be a continuous map equipped with a fitting flow $\Phi$ in a neighborhood of $x\in M$. For $t$ small enough the sphere $S_{t}:[q]\in\mathbb{S}u^{*}\mathcal{E}_{|x}\mapsto p\circ\Phi_{t}([q])\in M$ is homotopic to a generator of the homology of $U\setminus\{x\}$ for any open neighborhood $U$ of $x$ in $M$ that is diffeomorphic to $\mathbb{R}^{d}$.
@@
Let $P=u(x)$. Let $p_{0}\in P^{\circ}\subset S^{2}V$ be a positive tensor, which exists since the pencil $P$ is assumed to be in $\operatorname{Gr}^{{\operatorname{mix}}}_{d}(\mathcal{Q})$. Let $P^{\prime}\in\operatorname{Gr}_{d}(S^{2}V)$ be a supplement of $P^{\circ}$. Since $u$ is continuous for all $y$ close enough to $x$ the vector subspace $u^{\circ}(y)$ is transverse to $P^{\prime}$, therefore there exist a unique vector $\phi(y)\in P^{\prime}$ such that:
$p_{0}+\phi(y)\in(p_{0}+P^{\prime})\cap u^{\circ}(y)\subset S^{2}V^{>0}.$
@@
Let $[q]\in\mathbb{S}u^{*}\mathcal{E}_{x}$. For all $t$ such that $p\circ\Phi_{t}(q)\in U$, the linear form $\pi\left(\Phi_{t}(q)\right)\in\mathcal{Q}=S^{2}V^{*}$ vanishes on $\phi\left(p\circ\Phi_{t}(q)\right)\in S^{2}V$ since this point belongs to $u^{\circ}(p\circ\Phi_{t}(q))$. Moreover $\pi\left(\Phi_{t}(q)\right)-\pi(q)$ is a positive bilinear form since $\Phi$ is a fitting flow.
In particular $\pi(q)\in\mathcal{Q}=S^{2}V^{*}$ is always negative on $\phi\left(p\circ\Phi_{t}(q)\right)\in S^{2}V$. Hence for $t$ small enough $[\phi\circ S_{t}]:\mathbb{S}u^{*}\mathcal{E}_{x}\to\mathbb{S}P^{\prime}$ has the same degree as $[\pi]:\mathbb{S}u^{*}\mathcal{E}_{x}\to\mathbb{S}P^{\prime*}\simeq\mathbb{S}P$ which associates to $[q]\in\mathbb{S}u^{*}\mathcal{E}_{x}$ the class $[\pi(q)]\in\mathbb{S}P$. The map $[\pi]$ is a diffeomorphism, so in particular $1=|\deg([\pi])|=|\deg(\phi)\deg(S_{t})|$. Hence $S_{t}$ is a generator of the homotopy group of $U\setminus\{x\}$.
@@
When $M=\widetilde{N}$ where $N$ is compact and $\Gamma=\pi_{1}(M)$, the quotient $\mathbb{S}u^{*}\mathcal{E}/\rho(\Gamma)$ is compact. Hence any Riemannian metric on $\mathbb{S}u^{*}\mathcal{E}/\rho(\Gamma)$ is quasi-isometric to $\Gamma$ via any orbit map. When a fitting flow exists, the flow lines project to quasi-geodesics in $\widetilde{N}$.
@@
In this proof we will use the Hilbert distance $d_{H}$ on the properly convex domain $\mathbb{P}(S^{2}V^{>0})$. It is defined which is defined using the cross ratio as $d_{H}([p_{1}],[p_{2}])=\log\left(\text{cr}\left([p^{-}],[p_{1}],[p_{2}],[p^{+}]\right)\right)$ where $[p_{1}],[p_{2}]\in\mathbb{P}(S^{2}V^{>0})$ and $[p^{-}]$, $[p^{+}]$ are the intersection of the projective line through $[p_{1}]$ and $[p_{2}]$ with the boundary of the domain $\mathbb{P}(S^{2}V^{>0})$. If the closure in $\mathbb{P}(S^{2}V)$ of two sets $A,B\subset\mathbb{P}(S^{2}V^{>0})$ are disjoint, then the two sets are at positive distance for the Hilbert distance. Indeed for every Riemannian metric $d_{\mathbb{S}}$ on the compact manifold $\mathbb{P}(S^{2}V)$, the Hilbert distance between any two points in $\mathbb{P}(S^{2}V^{>0})$ is bounded from below by some uniform multiple of $d_{\mathbb{S}}$.
@@
For all $[q]\in\mathbb{S}u^{*}\mathcal{E}$ we consider the convex $\mathbb{P}(\langle\pi(q)\rangle^{\circ}\cap S^{2}V^{>0})$.
We choose a continuous and $\rho$-equivariant map $s:\mathbb{S}u^{*}\mathcal{E}\to\mathbb{P}(S^{2}V^{>0})$ with the property that $s[q]\in\mathbb{P}(\langle\pi(q)\rangle^{\circ}\cap S^{2}V^{>0})$ for all $[q]\in\mathbb{S}u^{*}\mathcal{E}$.
@@
We fix a $\Gamma$-invariant Riemannian metric $g$ on $\mathbb{S}u^{*}\mathcal{E}$ with associated distance $d_{g}$. We set $C_{1}$ to be the supremum of $d_{H}\left(s([q_{1}]),s([q_{2}])\right)$ for all $[q_{1}],[q_{2}]\in\mathbb{S}u^{*}\mathcal{E}$ such that $d_{g}([q_{1}],[q_{2}])\leq 1$, which exists since $\Gamma$ acts cocompactly on $\widetilde{N}$. The following inequality follows for all $[q_{1}],[q_{2}]\in\mathbb{S}u^{*}\mathcal{E}$ from the triangular inequality:
$d_{H}\left(s([q_{1}]),s([q_{2}])\right)\leq C_{1}d_{g}([q_{1}],[q_{2}])+C_{1}.$
@@
Indeed for some integer $n\leq d_{g}([q_{1}],[q_{2}])+1$ one can find elements $x_{0}=[q_{1}],x_{1},\cdots,x_{n}=[q_{2}]$ in $\mathbb{S}u^{*}\mathcal{E}$ such that for all $1\leq i\leq n$ one has $d_{g}(x_{i-1},x_{i})\leq 1$. The triangular inequality for the Hilbert distance implies that:
$d_{H}\left(s([q_{1}]),s([q_{2}])\right)\leq nC_{1}.$
@@
We now set $C_{2}$ to be the supremum of $d_{g}\left(s([q_{1}]),s(\Phi_{t}([q]))\right)$ for all $[q_{1}]\in\mathbb{S}u^{*}\mathcal{E}$ and $0\leq t\leq 1$. Similarly we get the following inequality for all $[q]\in\mathbb{S}u^{*}\mathcal{E}$ and $t\geq 0$:
$d_{g}\left(s([q]),s(\Phi_{t}([q]))\right)\leq C_{2}t+C_{2}.$
@@
Let $K$ be a compact fundamental domain for the action of $\Gamma$ on $\mathbb{S}u^{*}\mathcal{E}$ and let $\epsilon$ be the infimum of the Hilbert distance for any $[q]\in K$ between $\mathbb{P}\left(\langle\pi(q)\rangle^{\circ}\cap S^{2}V^{>0}\right)$ and $\mathbb{P}\left(\langle\pi(\Phi_{1}(q))\rangle^{\circ}\cap S^{2}V^{>0}\right)$. Since the flow is fitting, the closures of these two sets in $\mathbb{P}(S^{2}V)$ are disjoint for any $[q]$, and hence their Hilbert distance is positive. Since $K$ is compact, the infimum $\epsilon$ is also positive.
@@
The Hilbert distance between $s(q)$ and $s(\Phi_{t}(q))$ for $t>0$ and $q\in\mathbb{S}u^{*}\mathcal{E}$ is greater than $\epsilon(t-1)$. Indeed for all integer $0\leq n\leq t$ the projective segment between $s(q)$ and $s(\Phi_{t}(q))$, which is a geodesic for the Hilbert distance, intersects $\mathbb{P}\left(\langle\pi(\Phi_{n}(q))\rangle^{\circ}\cap S^{2}V^{>0}\right)$ in exactly one point $x_{n}$. Moreover the Hilbert distance between $x_{n}$ and $x_{n+1}$ for $10\leq n\leq t-$ 1 is at least $\epsilon$.
Putting all of these inequalities together we get that for all $t\geq 0$ and $[q]\in\mathbb{S}u^{*}\mathcal{E}$:
$\frac{\epsilon}{C_{1}}(t-1)-1\leq d_{g}\left(s([q]),s(\Phi_{t}([q]))\right)\leq C_{2}(t+1)$
@@
We now check that flow lines exist between any pair of points. Let $x\in M$. Given $t\in\mathbb{R}$ we consider the $d$-sphere $S_{t}:q\in\mathbb{S}u^{*}\mathcal{E}_{|x}\mapsto\Phi_{t}(q)\in u^{*}\mathcal{E}$. Suppose that some $y\in M$, avoids the sphere $S_{t}$ for all $t>0$. Consider a curve $\eta$ between $x$ and $y$. The homological intersection between this segment and the spheres $S_{t}$ in $M\setminus\{x,y\}$ is constant, and is equal to zero for $t$ large enough since the spheres $S_{t}$ are then uniformly far from $x$. However for $t$ small enough, the homotopy class of $S_{t}$ is the one of any small sphere encircling $x$ by Lemma 4.3. This leads to a contradiction since such a sphere will have homological intersection equal to $1$ or $-1$ with the curve $\eta$. Hence there exists a flow line joining any pair of points.
@@
Let $u:\widetilde{N}\to\operatorname{Gr}^{{\operatorname{mix}}}_{d}(\mathcal{Q})$ be a $\rho$-equivariant continuous map that admits an equivariant fitting flow, then it is an embedding. In particular it is a globally fitting map. Moreover $\rho:\Gamma\to G$ is a quasi-isometric embedding.

### v2

@@ 3.1. Pencils of quadrics
A vector $\mathrm{v}\in T_{P}\operatorname{Gr}_{d}(\mathcal{Q})$ is fitting if and only if, for every $\mathcal{C}^{1}$ curve $\gamma:[0,1]\to\operatorname{Gr}_{d}(\mathcal{Q})$ with $\gamma_{0}=P$ and $\gamma^{\prime}_{0}=\mathrm{v}$, the pair $(\gamma_{t},\gamma_{0})$ is fitting for all $t>0$ small enough.
@@
Let us fix a complement $H$ of $P^{\circ}$ in $S^{2}V$. Let $N=\dim(\mathcal{Q})=\frac{n(n+1)}{2}$. We identify a neighborhood of $P^{\circ}\subset\operatorname{Gr}_{N-d}(\mathcal{Q})$ with $\operatorname{Hom}(P^{\circ},H)$. If there exist a non-zero element $s\in\operatorname{Ker}(\mathrm{v}^{\circ})\cap S^{2}V^{\geq 0}$, one can consider the curve where $\gamma_{t}$ corresponds to
@@
The non-zero element $s\in S^{2}V^{\geq 0}$ belongs to $P^{\circ}=\gamma_{0}^{\circ}$ and $\gamma_{t}^{\circ}$ hence the pair $(\gamma_{0},\gamma_{t})$ is not fitting for any $t>0$.
In general the element corresponding to $\gamma_{t}$ is equal for $t$ close to 0 to the element of $\operatorname{Gr}_{N-d}(\mathcal{Q})$ which is the graph of the map $P^{\circ}\to H$:
@@
In this case the distance between $\mathbb{P}(\gamma_{0}^{\circ}\cap S^{2}V^{\geq 0})$ and $\mathbb{P}(\gamma_{t}^{\circ}\cap S^{2}V^{\geq 0})$ grows at least linearly in $t$ for $t$ close to 0. ∎
@@
Because of the second part of Definition 3.6 a vector is well-fitting if and only if it belongs to $C_{[q]}$ for some $q]\in P$. We just check that the sets $C_{[q]}$ are indeed open convex cones.
@@
Note here that for any linear subspace $A\subset V$ the projective codimension of $P(A)$ in $\mathbb{P}(V)$ is equal to the codimension of $A$ in $V$.
@@
Consider the tautological rank $d$ vector bundle $\operatorname{pr}:\mathcal{E}\to\operatorname{Gr}_{d}(\mathcal{Q})$. The fiber at $P\in\operatorname{Gr}_{d}(\mathcal{Q})$ of bundle $\mathcal{E}$ is identified with the vector subspace $P\subset\mathcal{Q}$. Since all the fibers are naturally identified with subsets of $\mathcal{Q}$, there is a tautological projection $\pi:\mathcal{E}\to\mathcal{Q}$ such that for $(P,q)$ in the tautological bundle, i.e. $P\in\operatorname{Gr}_{d}(\mathcal{Q})$ and $q\in P$ with $q\in P$, $\pi(P,q)=q$. We still denote by $\pi:u^{*}\mathcal{E}\to\mathcal{Q}$ the corresponding projection with a slight abuse of notations.
Note that that an element in $T\mathcal{Q}$ is a pair $(q,\dot{q})\in T_{q}\mathcal{Q}$ and we say that it is positive if the symmetric bilinear form $\dot{q}\in\mathcal{Q}$ is positive.
@@
    for all $v\in T_{x_{0}}M$ there exist $\mathrm{w}\in T\mathcal{E}$ such that $\mathrm{d}\operatorname{pr}(\mathrm{w})=\mathrm{d}u(v)$ and $\mathrm{d}\pi(\mathrm{w})\in T\mathcal{Q}$ is positive.
@@
A pair $(P,q_{0})\in\operatorname{Gr}_{d}(\mathcal{Q})\times\mathcal{Q}$ correspond to an element in $\mathcal{E}$ if and only if $q_{0}\in P$. A pair $\left(\mathrm{v},\dot{q}\right)\in T_{P}\operatorname{Gr}_{d}(\mathcal{Q})\times T_{q}\mathcal{Q}$ can be written as $\left(\mathrm{d}\operatorname{pr}(\mathrm{w}),\mathrm{d}\pi(\mathrm{w})\right)$ for some $\mathrm{w}\in T\mathcal{E}$ if and only if $\dot{q}$ belongs to the class defined by $\mathrm{v}(q_{0})$.
@@
The pullback $u^{*}\mathcal{E}$ of the tautological bundle $\operatorname{pr}:\mathcal{E}\to\operatorname{Gr}_{d}(\mathcal{Q})$ defines a rank $d$ vector bundle over $M$. We define the sphere bundle $\mathbb{S}u^{*}\mathcal{E}$ as the quotient of the vector bundle $u^{*}\mathcal{E}$ by the action of positive scalars.
Recall that $M$ has dimension $d$. We consider flows on $\mathbb{S}u^{*}\mathcal{E}$ so that some form of contraction occurs along the flow lines. We denote also by $\operatorname{pr}$ the bundle maps $u^{*}\mathcal{E}\to M$, $\mathbb{S}u^{*}\mathcal{E}\to M$, with a slight abuse of notations.
The fiber at $P\in\operatorname{Gr}_{d}(\mathcal{Q})$ of the bundle $\mathcal{E}$ is identified with the vector subspace $P\subset\mathcal{Q}$. Recall that there is a natural projection $\pi:\mathcal{E}\to\mathcal{Q}$ defined by $\pi(P,q)=q$. We still denote by $\pi:u^{*}\mathcal{E}\to\mathcal{Q}$ the corresponding projection with a slight abuse of notations.
@@
A fitting flow for a continuous map $u:M\to\operatorname{Gr}_{d}(\mathcal{Q})$ is a continuous flow $(\Phi_{t})_{t\in\mathbb{R}}$ on $\mathbb{S}u^{*}\mathcal{E}$ such that
one can choose representatives $q$ of $[q]$ and $q^{\prime}$ of $q^{\prime}]=\Phi_{t}([q])$ such that $\pi(q^{\prime})-\pi(q)\in\mathcal{Q}$ is positive.
@@
Assume by contradiction that for some $t_{0}>0$ and $q\in\mathbb{S}u^{*}\mathcal{E}$ one has $u(x)=\operatorname{pr}(q)=\operatorname{pr}(\Phi_{t_{0}}(q))$. The fact that the flow is fitting implies that for some $\lambda>0$, $\lambda\pi(\Phi_{t_{0}}(q))-\pi(q)\in\mathcal{Q}$ is positive. This positive quadric belongs to $u(x)$, contradicting the fact that $u(x)\in\operatorname{Gr}_{d}^{{\operatorname{mix}}}(\mathcal{Q})$ by Proposition 3.2.
@@
Let $u:M\to\operatorname{Gr}^{{\operatorname{mix}}}_{d}(\mathcal{Q})$ be a continuous map equipped with a fitting flow $\Phi$ in a neighborhood of $x\in M$. For $t>0$ small enough the sphere $S_{t}:[q]\in\mathbb{S}u^{*}\mathcal{E}_{|x}\mapsto p\circ\Phi_{t}([q])\in M$ is homotopic to a generator of the homology of $U\setminus\{x\}$ for any open neighborhood $U$ of $x$ in $M$ that is diffeomorphic to $\mathbb{R}^{d}$.
@@
Let $P=u(x)$. Let $s_{0}\in P^{\circ}\subset S^{2}V$ be a positive tensor, which exists since the pencil $P$ is assumed to be in $\operatorname{Gr}^{{\operatorname{mix}}}_{d}(\mathcal{Q})$. Let $P^{\prime}\in\operatorname{Gr}_{d}(S^{2}V)$ be a supplement of $P^{\circ}$. Since $u$ is continuous for all $y$ close enough to $x$ the vector subspace $u^{\circ}(y)$ is transverse to $P^{\prime}$, therefore there exist a unique vector $\phi(y)\in P^{\prime}$ such that:
$s_{0}+\phi(y)\in(s_{0}+P^{\prime})\cap u^{\circ}(y)\subset S^{2}V^{>0}.$
@@
Let $q]\in\mathbb{S}u^{*}\mathcal{E}_{x}$. For all $t$ such that $s\circ\Phi_{t}(q)\in U$, the linear form $\pi\left(\Phi_{t}(q)\right)\in\mathcal{Q}=S^{2}V^{*}$ vanishes on $\phi\left(s\circ\Phi_{t}(q)\right)\in S^{2}V$ since this point belongs to $u^{\circ}(\operatorname{pr}\circ\Phi_{t}(q))$. Moreover $\pi\left(\Phi_{t}(q)\right)-\pi(q)$ is a positive bilinear form since $\Phi$ is a fitting flow.
In particular $\pi(q)\in\mathcal{Q}=S^{2}V^{*}$ is always negative on $\phi\left(\operatorname{pr}\circ\Phi_{t}(q)\right)\in S^{2}V$. Hence for $t$ small enough $\phi\circ S_{t}]:\mathbb{S}u^{*}\mathcal{E}_{x}\to\mathbb{S}P^{\prime}$ has the same degree as $\pi]:\mathbb{S}u^{*}\mathcal{E}_{x}\to\mathbb{S}P^{\prime*}\simeq\mathbb{S}P$ which associates to $q]\in\mathbb{S}u^{*}\mathcal{E}_{x}$ the class $\pi(q)]\in\mathbb{S}P$. The map $\pi]$ is a diffeomorphism, so in particular $1=|\deg([\pi])|=|\deg(\phi)\deg(S_{t})|$. Hence $S_{t}$ is a generator of the homotopy group of $U\setminus\{x\}$.
@@
Let $M=\widetilde{N}$ where $N$ is compact and $\Gamma=\pi_{1}(N)$. Let $\rho:\Gamma\to\operatorname{SL}(V)$ be a representation. Recall that a $(C,D)$-quasi geodesic in a metric space $(X,d)$ is a curve $\gamma:\mathbb{R}\to X$ such that for all $t_{1},t_{2}\in\mathbb{R}$:
$1/C|t_{1}-t_{2}|-D\leq d(\gamma(t_{1}),\gamma(t_{2}))\leq C|t_{1}-t_{2}|+D.$
@@
In the previous statement one can consider any fixed Riemannian metric on the compact space $\mathbb{S}u^{*}\mathcal{E}/\rho(\Gamma)$.
@@
In this proof we will use the Hilbert distance $d_{H}$ on the properly convex domain $\mathbb{P}(S^{2}V^{>0})$. It is defined which is defined using the cross ratio as $d_{H}([s_{1}],[s_{2}])=\log\left(\text{cr}\left([s^{-}],[s_{1}],[s_{2}],[s^{+}]\right)\right)$ where $s_{1}],[s_{2}]\in\mathbb{P}(S^{2}V^{>0})$ and $^{-}], [s+][s^{+}$] are the intersection of the projective line through $s_{1}]$ and $s_{2}]$ with the boundary of the domain $\mathbb{P}(S^{2}V^{>0})$. If the closure in $\mathbb{P}(S^{2}V)$ of two sets $A,B\subset\mathbb{P}(S^{2}V^{>0})$ are disjoint, then the two sets are at positive distance for the Hilbert distance. Indeed for every Riemannian metric $d_{\mathbb{S}}$ on the compact manifold $\mathbb{P}(S^{2}V)$, the Hilbert distance between any two points in $\mathbb{P}(S^{2}V^{>0})$ is bounded from below by some uniform multiple of $d_{\mathbb{S}}$.
@@
For all $q]\in\mathbb{S}u^{*}\mathcal{E}$ we consider the convex $\mathbb{P}(\langle\pi(q)\rangle^{\circ}\cap S^{2}V^{>0})$.
We choose a continuous and $\rho$-equivariant map $\sigma:\mathbb{S}u^{*}\mathcal{E}\to\mathbb{P}(S^{2}V^{>0})$ with the property that $\sigma([q])\in\mathbb{P}(\langle\pi(q)\rangle^{\circ}\cap S^{2}V^{>0})$ for all $q]\in\mathbb{S}u^{*}\mathcal{E}$. One can always find such a choice, as it amonts to finding a section of an open ball bundle over a compact set. Indeed $\mathbb{P}(\langle\pi(q)\rangle^{\circ}\cap S^{2}V^{>0})$ is a non-empty convex set of codimension $1$ that forms an open ball for all $q]\in\mathbb{S}u^{*}\mathcal{E}$.
@@
We fix a $\Gamma$-invariant Riemannian metric $g$ on $\mathbb{S}u^{*}\mathcal{E}$ with associated distance $d_{g}$. We set $C_{1}$ to be the supremum of $d_{H}\left(\sigma([q_{1}]),\sigma([q_{2}])\right)$ for all $q_{1}],[q_{2}]\in\mathbb{S}u^{*}\mathcal{E}$ such that $d_{g}([q_{1}],[q_{2}])\leq 1$, which exists since $\Gamma$ acts cocompactly on $\widetilde{N}$. The following inequality follows for all $q_{1}],[q_{2}]\in\mathbb{S}u^{*}\mathcal{E}$ from the triangular inequality:
$d_{H}\left(\sigma([q_{1}]),\sigma([q_{2}])\right)\leq C_{1}d_{g}([q_{1}],[q_{2}])+C_{1}.$
@@
Let $n$ be the unique integer $d_{g}([q_{1}],[q_{2}])\leq n<d_{g}([q_{1}],[q_{2}])+1$. One can find elements $x_{0}=[q_{1}],x_{1},\cdots,x_{n}=[q_{2}]$ in $\mathbb{S}u^{*}\mathcal{E}$ such that for all $1\leq i\leq n$ one has $d_{g}(x_{i-1},x_{i})\leq 1$. The triangular inequality for the Hilbert distance implies that:
$d_{H}\left(\sigma([q_{1}]),\sigma([q_{2}])\right)\leq nC_{1}.$
@@
We now set $C_{2}$ to be the supremum of $d_{H}\left(\sigma([q_{1}]),\sigma(\Phi_{t}([q]))\right)$ for all $q_{1}]\in\mathbb{S}u^{*}\mathcal{E}$ and $0\leq t\leq 1$. Similarly we get the following inequality for all $q]\in\mathbb{S}u^{*}\mathcal{E}$ and $t\geq 0$:
$d_{H}\left(\sigma([q]),\sigma(\Phi_{t}([q]))\right)\leq C_{2}t+C_{2}.$
@@
Let $K$ be a compact fundamental domain for the action of $\Gamma$ on $\mathbb{S}u^{*}\mathcal{E}$ and let $\epsilon$ be the infimum of the Hilbert distance for any $q]\in K$ between $\mathbb{P}\left(\langle\pi(q)\rangle^{\circ}\cap S^{2}V^{>0}\right)$ and $\mathbb{P}\left(\langle\pi(\Phi_{1}(q))\rangle^{\circ}\cap S^{2}V^{>0}\right)$. Since the flow is fitting, the closures of these two sets in $\mathbb{P}(S^{2}V)$ are disjoint for any $[q]$, and hence their Hilbert distance is positive. Since $K$ is compact, the infimum $\epsilon$ is also positive.
@@
The Hilbert distance between $\sigma(q)$ and $\sigma(\Phi_{t}(q))$ for $t>0$ and $q\in\mathbb{S}u^{*}\mathcal{E}$ is greater than $\epsilon(t-1)$. Indeed for all integer $0\leq n\leq t$ the projective segment between $\sigma(q)$ and $\sigma(\Phi_{t}(q))$, which is a geodesic for the Hilbert distance, intersects $\mathbb{P}\left(\langle\pi(\Phi_{n}(q))\rangle^{\circ}\cap S^{2}V^{>0}\right)$ in exactly one point $x_{n}$. Hence the Hilbert distance between $x_{n}$ and $x_{n+1}$ for $0\leq n\leq t-1$ is at least $\epsilon$.
Putting all of these inequalities together we get that for all $t\geq 0$ and $q]\in\mathbb{S}u^{*}\mathcal{E}$:
$\frac{\epsilon}{C_{1}}(t-1)-1\leq d_{g}\left(\sigma([q]),\sigma(\Phi_{t}([q]))\right)\leq C_{2}(t+1)$
@@
We now check that flow lines exist between any pair of points. Let $x\in M$. Given $t\in\mathbb{R}$ we consider the $d$-sphere $S_{t}:q\in\mathbb{S}u^{*}\mathcal{E}_{|x}\mapsto\Phi_{t}(q)\in u^{*}\mathcal{E}$. Suppose that some $y\in M$, avoids the sphere $S_{t}$ for all $t>0$. Consider a curve $\eta$ between $x$ and $y$. The homological intersection between this segment and the spheres $S_{t}$ in $M\setminus\{x,y\}$ is constant, and is equal to zero for $t$ large enough since the spheres $S_{t}$ are then uniformly far from $x$. However for $t$ small enough, the homotopy class of $S_{t}$ is the one of any small sphere encircling $x$ by Lemma 4.3. This leads to a contradiction since such a sphere will have homological intersection equal to $1$ or 1-1 with the curve $\eta$. Hence there exists a flow line joining any pair of points.
@@
Let $u:\widetilde{N}\to\operatorname{Gr}^{{\operatorname{mix}}}_{d}(\mathcal{Q})$ be a $\rho$-equivariant continuous map that admits an equivariant fitting flow, then it is an embedding. In particular it is a globally fitting map. Moreover $\rho:\Gamma\to\operatorname{SL}(V)$ is a quasi-isometric embedding.